# Towards a Science of Scaling Agent Systems

### v1

@@ 6 Conclusion
This study quantifies scaling principles for agentic systems across 180 controlled experiments spanning three LLM families and four agentic benchmarks. We reveal that multi-agent performance exhibits an inverted-U relationship with coordination complexity, with benefits diminishing beyond moderate coordination levels. Domain complexity emerges as the strongest performance predictor ($\beta=-0.114$, $p<0.002$), reducing MAS advantage more substantially than architectural choices. Performance gains vary dramatically by task structure: +80.9% on Finance Agent versus -70.0% on PlanCraft, indicating that coordination benefits depend tightly on task decomposability. Under fixed computational budgets, turn count grows superlinearly with team size ($T\propto n^{1.724}$), constraining effective team sizes to 3–4 agents in practice. We derive a predictive relationship showing that MAS advantage depends on both model capability ($I$) and domain complexity ($D$), enabling practitioners to estimate expected gains before deployment. Cross-validation achieves $R^{2}=0.89$ on held-out data. Our findings suggest that multi-agent benefits depend critically on task structure rather than team size alone. Effective system design requires matching coordination topology to problem characteristics, rather than assuming uniform benefits from scaling agent count.
@@ Appendix A Model Intelligence Index
To quantify the capabilities of LLMs used in our study, we adopt while extending the Artificial Analysis Intelligence Index [^3] https://artificialanalysis.ai/evaluations/artificial-analysis-intelligence-index. This index provides one of the most comprehensive publicly available syntheses of model capabilities, combining performance across reasoning, knowledge, mathematics, coding, instruction following, long-context reasoning, and agentic workflow tasks. Its construction integrates ten evaluation suites (e.g., MMLU-Pro (wang2024mmlu), GPQA Diamond (rein2024gpqa), HLE (phan2025humanity), AIME 2025, SciCode (tian2024scicode), LiveCodeBench (jain2025livecodebench), IFBench (pyatkin2025generalizing), AA-LCR (artificialanalysis2025lcr), Terminal-Bench Hard, and $\tau^{2}$-Bench Telecom (barres2025t2bench)), with careful standardization, robust answer extraction, and model-agnostic prompting.
@@
For the models without publicly available results, we computed a reconstructed Intelligence Index following the equal-weighting formulation used in Intelligence Index v3.0. In cases where full reproduction was infeasible (e.g., specific agentic workflow tasks or unavailable context window limits), we report approximate estimates (denoted with $\sim$”) and discuss their limitations transparently. These reconstructed values should be interpreted as methodologically consistent but not officially certified estimates.
@@
Model Index AA-LCR HLE MMLU GPQA AIME LiveCode SciCode IFBench
[Uncaptioned image] GPT-5 66 78 27 87 85 85 54 37 94
[Uncaptioned image] GPT-5 mini 62 68 20 84 83 84 39 75 91
[Uncaptioned image] GPT-5 nano 49 74 28 78 68 79 37 68 84
[Uncaptioned image] Gemini-2.5 Pro 58 54 21 86 84 87 70 43 88
[Uncaptioned image] Gemini-2.5 Flash 52 64 13 84 79 71 41 52 78
[Uncaptioned image] Gemini-2.0 Flash 34 45 8 77 68 73 39 35 70
[Uncaptioned image] Claude 4.5 Sonnet 61 66 17 88 83 71 45 57 88
[Uncaptioned image] Claude 4.0 Sonnet 57 62 15 87 75 71 56 48 85
[Uncaptioned image] Claude 3.7 Sonnet 48 58 12 81 67 22 57 42 80
@@
Appendix B Domain Complexity
@@
B.1 Complexity Metric Construction
@@ Appendix C Datasets
Appendix D Implementation Details
D.1 Technical Infrastructure
@@ D.2 Agent Configuration
D.3 Prompt Compilation System
@@ D.4 Evaluation Methodology
D.5 Information Gain Computation
@@
Figure 6: Benchmark-specific scaling dynamics across LLM families. Performance curves across four benchmarks show best-performing multi-agent variants versus single-agent baselines by Intelligence Index. OpenAI and Google exhibit strong cooperative scaling in structured tasks (Finance Agent: $23.1\%$; Workbench: $20.8\%$; Cohen’s $d>1.2$). Anthropic models show diminished or negative returns in open-ended environments (PlanCraft: $35.0\%$ for uncoordinated variants; $d\approx 0.35$), where independent reasoning sometimes outperforms coordination. Cross-family variance decomposition ($R^{2}=0.89$) confirms that intrinsic communication alignment differences drive these divergent patterns.

### v2

@@ 6 Conclusion
[…]
@@ Appendix A Model Intelligence Index
To quantify the capabilities of LLMs used in our study, we adopt while extending the Artificial Analysis Intelligence Index [^3] https://artificialanalysis.ai/evaluations/artificial-analysis-intelligence-index. This index provides one of the most comprehensive publicly available syntheses of model capabilities, combining performance across reasoning, knowledge, mathematics, coding, instruction following, long-context reasoning, and agentic workflow tasks. Its construction integrates eight evaluation suites (e.g., MMLU-Pro (wang2024mmlu), GPQA Diamond (rein2024gpqa), HLE (phan2025humanity), AIME 2025, SciCode (tian2024scicode), LiveCodeBench (jain2025livecodebench), IFBench (pyatkin2025generalizing), AA-LCR (artificialanalysis2025lcr), Terminal-Bench Hard, and $\tau^{2}$-Bench Telecom (barres2025t2bench)), with careful standardization, robust answer extraction, and model-agnostic prompting.
@@
For the models without publicly available results, we computed a reconstructed Intelligence Index following the equal-weighting formulation used in Intelligence Index v3.0. In cases where full reproduction was infeasible (e.g., specific agentic workflow tasks or unavailable context window limits), we report approximate estimates (denoted with *) and discuss their limitations transparently. These reconstructed values should be interpreted as methodologically consistent but not officially certified estimates.
@@
Model Index AA-LCR HLE MMLU-Pro GPQA Diamond AIME 25 LiveCode SciCode IFBench
[Uncaptioned image] GPT-5.2 75 73 31 87 90 99 89 52 75
[Uncaptioned image] GPT-5 71 76 27 87 85 94 85 43 73
[Uncaptioned image] GPT-5 mini 68 68 20 84 91 84 84 39 75
[Uncaptioned image] GPT-5 nano 59 42 8 78 84 79 79 37 68
[Uncaptioned image] Gemini-2.5 Pro 65 66 21 86 84 88 80 43 49
[Uncaptioned image] Gemini-2.5 Flash 58 57 13 84 79 78 63 41 52
[Uncaptioned image] Gemini-2.0 Flash 47 45 8 77 68 73 39 35 30
[Uncaptioned image] Claude 4.5 Sonnet 55 66 7 88 83 37 71 43 43
[Uncaptioned image] Claude 4.0 Sonnet 47 62 5 87 75 21 56 38 35
[Uncaptioned image] Claude 3.7 Sonnet 42 58 2 81 67 12 57 32 30
@@
Table 7: Out-of-sample validation metrics for GPT-5.2 on BrowseComp-Plus.
Metric Value Status
Mean Absolute Error (MAE) 0.071 <0.10<0.10
Mean Absolute Percentage Error (MAPE) 15.8% Acceptable
Normalized MAE 0.045
Qualitative Findings Validated 4/5 Partial
Kendall’s $\tau$ (Ranking) 0.200 Weak
Predictions normalized to actual performance range.
See Table 8; one finding shows partial validation due to architecture convergence at high capability.
Table 8: Validation of key findings from Section 4 on GPT-5.2 (held-out model). Four of five scaling principles derived from the training set generalize to higher-capability models.
Finding GPT-5.2 Result Valid
Capability Ceiling: $P_{\text{SA}}\geq 0.45\Rightarrow$ diminishing MAS returns ($\hat{\beta}=-0.404$) $P_{\text{SA}}=0.45$; best MAS gain $6.7\%$
Independent MAS Degradation: Independent architecture underperforms SAS Independent =0.35=0.35 vs SAS =0.45=0.45 ($22.2\%$)
Optimal Architecture: Centralized/Decentralized excel on BrowseComp-Plus Best: Centralized & Decentralized (tied at 0.48)
Hybrid Overhead: 515% overhead degrades performance Hybrid =0.39=0.39 vs SAS =0.45=0.45 ($13.3\%$)
BrowseComp Pattern: Decentralized >> Centralized on dynamic web tasks Decentralized =0.48=0.48, Centralized =0.48=0.48 Partial
Main results show Decentralized (+9.2%) outperforming Centralized (+0.2%); GPT-5.2 shows convergence, suggesting architecture-specific advantages may diminish at higher capability levels.
Table 9: Out-of-sample validation on GPT-5.2 (Intelligence Index = 75). Left: prediction accuracy by architecture. Right: validation of theoretical findings. The scaling equation (Eq. 1), trained on Index 42–71, achieves MAE = 0.071 and validates four of five qualitative predictions when extrapolating to higher-capability models.
Architecture Pred. Actual Error
MAS-Decentralized 0.496 0.480 ++3.3%
MAS-Hybrid 0.413 0.390 ++5.9%
MAS-Independent 0.344 0.350 -1.7%
MAS-Centralized 0.391 0.480 -18.6%
SAS 0.673 0.450 ++49.5%
MAE (MAS only) 0.034
MAE (Overall) 0.071
Finding Valid
Capability ceiling ($P_{\text{SA}}\geq 45\%$)
Independent MAS Degradation
Optimal architecture (Cent./Decent.)
Hybrid overhead degradation
[…]
Total 4/5
SAS over-prediction due to linear extrapolation beyond training range; quadratic term is not significant ($p=0.509$).
† Main results show Decentralized (+9.2%) outperforming Centralized (+0.2%); GPT-5.2 shows convergence (both at 0.48), suggesting architecture-specific advantages may diminish at higher capability levels.
Appendix B Out-of-Sample Validation
To assess the generalizability of our scaling equation beyond the training distribution, we evaluate on GPT-5.2 (Intelligence Index = 75), released after our experimental campaign. This represents extrapolation: our training data spans Index 42–71, while GPT-5.2 exceeds this range by 4% (5.6% beyond maximum).
Table 7 summarizes validation metrics. The scaling equation achieves MAE = 0.071, meeting our <0.10<0.10 threshold for practical utility. Critically, four of five qualitative findings from Section 4 generalize to this higher-capability regime (Table 8): the capability ceiling effect persists ($P_{SA}=0.45$, best MAS gain = +6.7%, consistent with diminishing returns at high baselines), Independent MAS exhibits predicted degradation (-22.2%), and Centralized/Decentralized architectures achieve optimal performance. The partial validation concerns the Decentralized >> Centralized pattern: while the main results show a clear Decentralized advantage over SAS (+9.2% vs. +0.2%), GPT-5.2 shows convergence (both at 0.48), suggesting that architecture-specific advantages may decrease as the model capability increases.
Architecture-wise analysis (Table 9) reveals that MAS predictions are well-calibrated (average error = 7.4%), while SAS shows systematic over-prediction (+49.5%). This bias does not stem from the quadratic intelligence term, which is not significant after centering ($\hat{\beta}_{I^{2}}=0.007$, $p=0.509$). Instead, the over-prediction likely reflects limitations in linear extrapolation of the capability effect ($\hat{\beta}_{I}=0.171$) beyond the training range (Index 42–71 to 75), where diminishing returns may emerge that are not captured by the current functional form. Future work should explore saturation functions or piecewise-linear specifications for extreme capability levels. These results validate our core contribution: the scaling equation captures architectural trade-offs that generalize across capability levels, even when absolute calibration requires refinement for extrapolation.
Architecture Selection Accuracy.
We note an important caveat regarding architecture selection when the model is at at extreme capability level. The scaling equation predicts SAS as optimal for GPT-5.2 ($\hat{P}_{\text{SAS}}=0.673$), yet empirically MAS-Centralized and MAS-Decentralized achieve superior performance (0.48 vs. 0.45). This discrepancy does not arise from the quadratic intelligence term, which is not significant after mean-centering ($\hat{\beta}_{I^{2}}=0.007$, $p=0.509$). Rather, the over-prediction reflects the linear intelligence effect ($\hat{\beta}_{I}=0.171$) extrapolating beyond its training range, potentially encountering diminishing returns at higher capability levels not captured by the current specification. The 87% architecture selection accuracy reported in Subsection 4.3 applies within the training range (Intelligence Index 42–71); extrapolation to Index 75 reduces this to 0% for SAS prediction but maintains 75% accuracy for MAS architecture ranking (3/4 correct: Decentralized, Hybrid, Independent). This suggests that while absolute performance predictions require recalibration for frontier models, relative MAS architecture rankings remain reliable.
Appendix C Domain Complexity
@@
C.1 Complexity Metric Construction
@@ Appendix D Datasets
Appendix E Implementation Details
E.1 Technical Infrastructure
@@ E.2 Agent Configuration
E.3 Prompt Compilation System
@@ E.4 Evaluation Methodology
E.5 Information Gain Computation
@@
Figure 6: Benchmark-specific scaling dynamics across LLM families. Performance curves across four benchmarks show best-performing multi-agent variants versus single-agent baselines by Intelligence Index. OpenAI and Google exhibit strong cooperative scaling in structured tasks (Finance Agent: $23.1\%$; Workbench: $20.8\%$; Cohen’s $d>1.2$). Anthropic models show diminished or negative returns in open-ended environments (PlanCraft: $35.0\%$ for uncoordinated variants; $d\approx 0.35$), where independent reasoning sometimes outperforms coordination.